OpenStudy (anonymous):

How do I convert y-5=-2(x+1)^2 to intercept form? The standard form is y=-2x^2-4x+3!

OpenStudy (anonymous):

You might want to check my work over, but if you are trying to get the equation into slope intercept y= mx+b form then these are the steps. First, multiply the quadratic in the parentheses; y-5=-2(x^2+1). then multiply -2; y-5=-2x^2-2. Then add 5 over; y=2x^2+3. I will check this through, but this is my best for now

OpenStudy (anonymous):

oops! i meant intercept form as in y=(x-r1)(x-r2)

OpenStudy (anonymous):

I tried factoring, but you can't factor it. and I tried the quadratic equation but i'm getting -1+- 2sqrt10 which isn't right i think

OpenStudy (anonymous):

If it is not factorizable, then I think you can't convert that equation into Intercept Form..

OpenStudy (anonymous):

really? idk i know the intercepts are -2.581 an .581 though

OpenStudy (anonymous):

and*

OpenStudy (anonymous):

I mean you can't convert, but the person who knows how to convert, that person can convert. :P

OpenStudy (anonymous):

Okay, let me try if I am that person or not.. :)

OpenStudy (anonymous):

Add 5 both sides first to separate \(y\)...

OpenStudy (anonymous):

\(y = -2(x^2 + 1 + 2x) + 5 \implies y = -2x^2 - 4x+ 3\)

OpenStudy (anonymous):

Let us go with Quadratic Formula for a moment: \[x_{1,2} = \frac{4 \pm 2\sqrt{10}}{-4} \\ x_{1,2} = -1 \mp \frac{\sqrt{10}}{2}\]

OpenStudy (anonymous):

So, yes you are right, you got the two intercepts right..

OpenStudy (anonymous):

So: \(y = (x - 0.581)(x + 2.581)\)

OpenStudy (anonymous):

ohh i thought i was wrong haha okay thanks:)